OpenStudy (anonymous):

WILL GIVE MEDALS!!!!! The measure of the second angle of a triangle is twice the measure of the first angle. The measure of the third angle is 10° less than the measure of the first angle. Find the measure of the second angle.

OpenStudy (anonymous):

The answer choices: A. 47.5° B. 80° C. 37.5° D. 95°

OpenStudy (welshfella):

let first angle = x then second angle = 2x

OpenStudy (anonymous):

Okay what next?

OpenStudy (welshfella):

and third angle = x - 10

OpenStudy (anonymous):

Then?

OpenStudy (welshfella):

well the 3 angles add up to 180 so we can write an equation any ideas ?

OpenStudy (anonymous):

Nope that's why I'm asking for help. If I knew how to set it up I could most likely do it.

OpenStudy (welshfella):

x + 2x + x - 10 = 180 makes sense?

OpenStudy (welshfella):

the 3 angles are on the left and 180 on the right so now if you solve for x youll find the measure of the first angle second angle = 2 8 x

OpenStudy (welshfella):

* 2 * x

OpenStudy (anonymous):

Okay how do I set up that part.

OpenStudy (welshfella):

add up the first 3 terms x + 2x + x = ?

OpenStudy (anonymous):

?

OpenStudy (welshfella):

the angle x + twice the angle x + angle x = how many angle x's?

OpenStudy (anonymous):

x + 2x + x = 3?

OpenStudy (welshfella):

1 + 2 + 1 = how many

OpenStudy (anonymous):

4, sorry I forgot to count both of the Xs for 2x.

OpenStudy (welshfella):

right 1 + 2 + 1 = 5 in same way x +2x + x = 4x

OpenStudy (welshfella):

* lol thats 4 NOT 5!!

OpenStudy (welshfella):

so we have 4x - 10 = 180 add 10 to both sides 4x - 10 + 10 = 180 + 10 4x = 190 x = 190/4

OpenStudy (welshfella):

now divide 190 by 4 to get x then multiply the result by 2 to get the values of 2x which is the values of the second angle

OpenStudy (anonymous):

190 / 4 = 47.5 47.5 * 2 = 95

OpenStudy (welshfella):

correct

OpenStudy (anonymous):

That's the answer to the whole problem ain't it?

OpenStudy (anonymous):

Thanks

OpenStudy (welshfella):

yes yw